OpenStudy (anonymous):

Use an inequality to solve the problem. Be sure to show the inequality and all of your work. One side of a rectangle is 8 inches and the other side is x inches. What values of x will make the perimeter at least 46 inches?

OpenStudy (anonymous):

that was wrong \[16+2x\geq 46\] \[2x\geq 30\] \[x\geq 15\] looks better